OpenStudy (anonymous):

To prove cos 2a = 2 cos^2 a - 1 What do I do from cos^2 a -sin^2 a

OpenStudy (asnaseer):

use the identity:\[\cos^2(a)+\sin^2(a)=1\]

OpenStudy (mathstudent55):

\( \sin^2 a + \cos^2 a = 1\) \( \cos 2a = 2 \cos^2 a - 1 \) \( \cos 2a = 2 \cos^2 a - (\sin^2 a + \cos^2 a) \) \( \cos 2a = 2 \cos^2 a - \sin^2 a - \cos^2 a \) \( \cos 2a = \cos^2 a - \sin^2 a \)

OpenStudy (anonymous):

Ok so then I get to 1 - 2 sin^2 a got it thanks again!

OpenStudy (asnaseer):

yw :)

OpenStudy (asnaseer):

you can also get it in terms of cos(a)

OpenStudy (anonymous):

Thank you mathstudent too

OpenStudy (mathstudent55):

wlcm

OpenStudy (anonymous):

awesome both very helpful :) Trying to learn Precalc in a week >.< found trig was the hardest part lol

OpenStudy (asnaseer):

I'm sure you'll get the hang of it soon - practice makes perfect! :)

OpenStudy (anonymous):

Luckily I have a ton of textbooks to use, just need to see some parts explained and I usually pick it up

OpenStudy (asnaseer):

thats great - the important point here is that you have the right spirit to learn the subject matter - a rare quality that you should be proud of! :)

OpenStudy (anonymous):

Haha thanks :) it's because I got moved up to Calc instead of Precalc this year for some reason and I dont want to get dropped and be bored in class for another year with over 100% xD

OpenStudy (asnaseer):

:)